A high school graduating class is made up of 601 students. There are 91 more girls than boys. how many boys are in the class?

Answer by fcabanski(1391)

b=boys
g=girls=b+91
g+b=601 which means b+91+b=601
2b=510
b=255
g = 255+91=346
It's better to understand math so you can solve these problems yourself. One on one online tutoring is the answer ($30/hr). For faster solutions with guaranteed detailed answers, use personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com